the half-life of radon-222 is 3.824 day. when you are 97 years old, you come across a 60 atom sample of Radon-222. After what time (in days) will 15 atoms of the original sample remain?
1
Expert's answer
2020-09-29T06:54:28-0400


Given,


the half-life of radon-222, (t)1/2 = 3.824 day


There are a 60 atom in sample of Radon-222.


We know that half life of radioactive sample, (t)1/2 = ln2/k


Where k = decay constant of the radioactive element.




Putting the values given,


3.824 = ln2/k


Or, k = ln2/3.824


Or, k ≈ 0.1813 /day




According to the integrated decay equation,


ln(N°/N) = kt_____[A]


where, N° = number of radioactive atoms present in the sample at the beginning.


N = number of radioactive atoms present in the sample at the time = t


Putting the values in the above equation [A] we get.


ln(60/15) = 0.1813 × t


Or, ln4 = 0.1813 × t


Or, t = ln4/0.1813


Or, t = 7.646 (up to three significant decimal)


Hence after 7.646 days 15 atoms of radon-222 to remain in the sample